OpenStudy (annalee607):

You need to switch the front and back tires ever 10,000 miles. Choose one vehicle and keep track of how far you travel in that vehicle for a week. Calculate the average, daily number of rotations made by the tires. Use your data to predict the date that you will need to witch the front and back tires. My choice: CRV Tires circumference 25 inches Travels about 225 miles in one week

TheSmartOne (thesmartone):

We need to convert miles into inches.

OpenStudy (annalee607):

10,000 miles =633,6000,000 inches

TheSmartOne (thesmartone):

1 mile = 5280 feet 1 feet = 12 inches How many inches is 225 miles?

OpenStudy (annalee607):

99,000

TheSmartOne (thesmartone):

That is wrong. 225 * 5280 * 12 will give you the answer The first question asks: Calculate the average, daily number of rotations made by the tires. So you drive 225 miles in 7 days. We need to find how much you drive on average in 1 day. We then convert that to inches. And then we divide by the circumference of your tire to find out how many rotations were made by your tires.

TheSmartOne (thesmartone):

For the second question: Use your data to predict the date that you will need to witch the front and back tires. All we have to do is 10,000/225 That gives us how many weeks later you will need to switch your tires.

OpenStudy (annalee607):

225*5280*12=14,256,000

TheSmartOne (thesmartone):

And divide that by 7 to get how many inches you travel in 1 day. And then we need to divide it by the circumference of your tire.

TheSmartOne (thesmartone):

When you said that the tire's circumference is 25, did you actually mean that the diameter of the tire is 25 inches?

OpenStudy (annalee607):

No it said nothing about Diameter . It wanted the circumference of the tire. 14,256,000/7=2,036,571.4286 2,036,571.4286/25=81,462.857144

TheSmartOne (thesmartone):

 So did you find the circumference like this?

OpenStudy (annalee607):

no I just measured straight across and it came out to 25

TheSmartOne (thesmartone):

When you measure straight across, that would be the diameter :)

OpenStudy (annalee607):

so the circumference would be 78.5

TheSmartOne (thesmartone):

So the circumference is \( 25\pi\) and \(\pi \approx 3.14\) and so we get: \(\Large \frac{2,036,571.4286}{78.5} = ?\)

TheSmartOne (thesmartone):

Yes

TheSmartOne (thesmartone):

What you get when you divide that is how many rotations your tires did in one day.

OpenStudy (annalee607):

2,594.3621478

TheSmartOne (thesmartone):

You're off by one decimal point. It's 25943.5851

OpenStudy (annalee607):

okay

TheSmartOne (thesmartone):

For the second question: Use your data to predict the date that you will need to witch the front and back tires. All we have to do is 10,000/225 That gives us how many weeks later you will need to switch your tires.

OpenStudy (annalee607):

44.444444444

TheSmartOne (thesmartone):

and them multiply that by 7 to find how many days and then you add the number of days to the current day to find out when you have to replace the tires

TheSmartOne (thesmartone):

We can use this to add the number of days http://www.timeanddate.com/date/dateadd.html So how many days are in 44.44 weeks? :)

OpenStudy (annalee607):

311.11

TheSmartOne (thesmartone):

http://www.timeanddate.com/date/dateadded.html?m1=5&d1=3&y1=2016&type=add&ay=&am=&aw=&ad=311&rec= and so we have to change the car tires on March 10th, 2017 That is considering we are using new tires starting today. But let's say you started using it last week since you had to see how many miles you drive on average. That would be 7 less days, which is March 3rd, 2017

OpenStudy (annalee607):

Thank you so much!(:

TheSmartOne (thesmartone):

And btw, nice car choice! CRV gets good miles per gallon. :)

TheSmartOne (thesmartone):

Anytime, Anna! (:

TheSmartOne (thesmartone):

And thank you for asking a qualified helper question!